OpenStudy (anonymous):

I know the equation for limits, but how do I find a limit for the measurement of one interior angle in a regular polygon?

OpenStudy (danjs):

the total of all the angles of a N sided polygon is = (N-2)*180

OpenStudy (danjs):

So one of the interior angles will be?

OpenStudy (anonymous):

What do you mean one of the interior angles will be?

OpenStudy (anonymous):

I'm not exactly sure what you mean

OpenStudy (danjs):

one interior angle of a polygon will be \[\frac{ (N-2)*180 }{ N }\]

OpenStudy (danjs):

Where N is the number of sides

OpenStudy (anonymous):

right my equation my teacher gave me is \[\frac{y= 180x-360 }{ x } \]

OpenStudy (anonymous):

I'm using a graph for this equation

OpenStudy (danjs):

ok, so what is the question?

OpenStudy (anonymous):

How do I use this equation to find the limit of the measure of one interior angle in a regular polygon?

OpenStudy (anonymous):

Does it matter what I use for x?

OpenStudy (danjs):

You want the limit as the number of sides goes to infinity?

OpenStudy (anonymous):

Yes, exactly.

OpenStudy (danjs):

right, well think about it, if you keep adding more and more sides, each angle is going to approach a straight line, but not quite

OpenStudy (anonymous):

Right...

OpenStudy (danjs):

so this is the equation

OpenStudy (anonymous):

Is there an actual number that can be achieved?

OpenStudy (danjs):

\[\lim_{x \rightarrow \infty} [ 180 - \frac{ 360 }{ x }]\]

OpenStudy (danjs):

As x gets very large, the fraction 360/x gets very small and approaches zero

OpenStudy (danjs):

The limit would be 180

OpenStudy (danjs):

Graph: y = 180 - 360/x look at what y approaches as you look at huge x values.

OpenStudy (anonymous):

But how does y get to 180 if it keeps getting smaller and smaller?

OpenStudy (anonymous):

Is there a specific number that can be achieved for y that I can physically write down?